### v1

# Cyber Defense Reinvented: Large Language Models as Threat Intelligence Copilots

Abstract
The exponential growth of cyber threat knowledge, exemplified by the expansion of databases such as MITRE-CVE and NVD, poses significant challenges for cyber threat analysis. Security professionals are increasingly burdened by the sheer volume and complexity of information, creating an urgent need for effective tools to navigate, synthesize, and act on large-scale data to counter evolving threats proactively. However, conventional threat intelligence tools often fail to scale with the dynamic nature of this data and lack the adaptability to support diverse threat intelligence tasks.
In this work, we introduce CyLens, a cyber threat intelligence copilot powered by large language models (LLMs). CyLens is designed to assist security professionals throughout the entire threat management lifecycle, supporting threat attribution, contextualization, detection, correlation, prioritization, and remediation. To ensure domain expertise, CyLens integrates knowledge from 271,570 threat reports into its model parameters and incorporates six specialized NLP modules to enhance reasoning capabilities. Furthermore, CyLens can be customized to meet the unique needs of different organizations, underscoring its adaptability. Through extensive evaluations, we demonstrate that CyLens consistently outperforms industry-leading LLMs and state-of-the-art cybersecurity agents. By detailing its design, development, and evaluation, this work provides a blueprint for leveraging LLMs to address complex, data-intensive cybersecurity challenges.
1 Introduction
We are living in an era of rapid digital transformation, where technological advancements are closely linked to the growing prevalence of cyber threats. In recent years, the landscape of cyber threats has evolved dramatically, marked by a 25% annual increase in reported CVEs [40]. The increase in CVEs can be attributed to several factors, especially the growing complexity of IT systems [76], the widespread adoption of open-source software [22, 79], and the accelerated pace of development cycles [83]. Together, these factors facilitate the exploitation of cyber threats in modern software ecosystems.
To enable efficiently analyze and counteract these threats, Cyber Threat Intelligence (CTI) [23] has emerged as a crucial strategy. CTI provides actionable insights derived from the analysis of threat data and empowers security teams to anticipate, identify, and mitigate cyber risks effectively [89, 110, 8].
However, CTI faces critical challenges in scalability and adaptability. First, the overwhelming volume of threat data creates significant bottlenecks [1]. Security teams are often inefficiently processing and correlating this massive influx of security knowledge, limiting their ability to respond to emerging cyber threats. Second, integrating CTI into existing cybersecurity platforms often requires extensive customization, which poses a significant barrier for small and medium-sized organizations who lack resources to deploy CTI effectively. To address these challenges, it is essential to propose advanced CTI solutions capable of scalably collecting and analyzing extensive cybersecurity knowledge while adaptively support customization tailored to diverse organizational needs.
Figure 1: (a) The lifecycle of threat management. Each task consists of multiple analytical targets, as detailed in Table 1. (b) The development methodology of CyLens, as detailed in section 3 and section 4.
Large language models (LLMs) have recently demonstrated exceptional capabilities, driven by their scalability to massive knowledge through pre-training and their adaptability for organization-specific applications via fine-tuning. Beyond these strengths, LLMs can be further augmented with task-oriented prompts or external modules, enabling more advanced analyses that require topic modeling [96], entity and relation recognition [97, 37], and retrieval-augmented generation (RAG) [33, 10]. Hence, LLMs emerge as a promising solution for assisting humans in specialized fields such as bug fixing [72, 94] and fuzz testing [38, 104, 24]. Consequently, we raise a critical question: how can LLMs be systematically developed as copilots to support CTI activities?
This work. To address this question, we propose CyLens, an LLM-powered copilot system designed to enhance CTI. CyLens consists of several sizes—1B, 8B, and 70B—to accommodate diverse needs and computational constraints. We characterize CyLens by the following aspects:
Lifecycle-Oriented Threat Management – As illustrated in Figure 1-(a), CyLens supports six tasks that cover the entire lifecycle of threat management. To analyze the properties of threats, CyLens enables ❶ attribution by identifying threat actors and gathering technical evidence, such as TTPs and exploitation paths, to trace the origin of a threat. It also facilitates ❷ contextualization by enriching metadata—such as affected systems—to provide a more comprehensive understanding of the threat landscape. Next, ❸ detection involves specifying available identifiers, such as CVE [14], while ❹ correlation enables CyLens to uncover adversarial networks by identifying relationships between the current threat and others. CyLens then drives ❺ prioritization by quantifying the severity and predicting the likelihood of exploitation. Building on this analytical foundation, ❻ remediation effectively counteracts threats and limits their broader impact.
Collective Expertise. As shown in Figure 1-(b), CyLens manages web-scale CTI knowledge through a combination of development methodologies. Our approach is motivated by the fact that cyber threats co-exist as an interdependent community, where attack vectors, infrastructure, and exploitations are deeply interconnected [46, 6, 93, 12, 103]. Effectively addressing these complex relationships requires a vast, integrated set of expertise rather than isolated knowledge.
To achieve this, CyLens first undergoes pre-training on a diverse set of databases—MITRE-CVE [14], CWE [16], CAPEC [17], ATT&CK [18], NVD [59], Exploit-DB [62], and correlated third-party repositories (e.g., [39, 5])—to incorporate extensive CTI knowledge from 271,570 threat reports. We introduce a curriculum pre-training approach, where “curriculum” data is designed to progressively embed CTI expertise into CyLens. Following pre-training, CyLens undergoes fine-tuning using task-oriented datasets addressing critical CTI tasks (❶-❻). To enhance its analytical depth, we employ a cascading reasoning process, allowing CyLens to systematically build awareness of dependencies within CTI analysis. Additionally, we integrate specialized design modules to reinforce CyLens ’s ability to conduct rigorous reasoning, escalating it as a human-usable CTI copilot.
Operational Proficiency. Our experiments demonstrate CyLens’s effectiveness across all above tasks under two categories of cyber threats: (i) Historical Threats—previously recorded threats that security analysts may not be fully aware of, and (ii) Zero-Day Threats—newly emerging threats that have not yet been documented. Across both threat categories, CyLens consistently outperforms SOTA models and industry-leading LLMs (e.g., ChatGPT and Gemini), underscoring its versatility in handling diverse cybersecurity challenges while excelling in proactive threat hunting within the ever-evolving cyberattack landscape.
Adaptive to Customization. CyLens also demonstrates adaptability in addressing organization-specific threats related to particular products, vendors, or threat groups. As shown in our evaluations in section 5.6, CyLens consistently delivers superior and balanced performance across diverse organizational requirements. This adaptability is particularly valuable when an organization’s threat-related data volume is insufficient, enabling effective CTI analysis even in data-scarce scenarios.
Contributions. We summarize following contributions:
    • LLM-as-CTI-Copilot: We propose CyLens, a scalable and adaptive system for CTI practice. CyLens is designed to accommodate diverse needs, making it well-suited for both research and industry applications.
    • Empirical Insights: Through extensive evaluations, we not only demonstrate CyLens’s effectiveness, but also reveal the varying performance of SOTA and industry-leading LLMs, highlighting the inherent complexity of CTI analysis and difficulty of achieving consistent expertise.
    • Blueprint Perspective: The design and development of CyLens serve as a blueprint for integrating LLM copilots into cybersecurity practice. This endeavor not only advances the capabilities of CTI systems but also provides insights that could drive the development of expert-level AI across other high-stakes domains [^1] Codes and datasets are organized at https://anonymous.4open.science/r/security-agent-4BB7/.
2 Background and Related Work
2.1 Cyber Threat Ingelligence (CTI)
AI-for-Security has been a long-standing research area, with extensive studies covering a broad range of cybersecurity topics, including anomaly detection [109, 68], malware classification[47, 69], and intrusion detection[48, 31]. Among these, CTI has emerged as a systematic approach to threat management, enabling organizations to collect, analyze, and operationalize intelligence on cyber threats [20, 95, 89].
Although there is no universal specification for CTI practices, we categorize CTI into six critical threat management tasks based on extensive literature [82, 32, 86, 54, 42, 75, 89, 71, 114, 43] and the MITRE initiative [3], a foundational effort in the field. As illustrated in Figure 1-(a), the lifecycle of threat intelligence begins with the extraction and analysis of threat properties (❶ & ❷) [42, 105, 75], progresses to the identification of adversary community (❸ & ❹) [32, 86, 54, 114], and culminates in the development of effective countermeasures (❺ & ❻) [89, 71, 43]. In this work, we set the scope of the CTI tasks as follows:
    • Attribution–Identifying the origin of cyber threats by profiling threat actors, analyzing tactics, techniques, and procedures (TTPs), and tracing exploitation paths to link cyber incidents to specific adversaries or groups. Attribution helps organizations understand attacker intent, capabilities, and historical attack patterns [105, 42].
    • Contextualization–Enriching raw threat intelligence with relevant metadata, such as affected systems, attack infrastructures, and potential impacts. Contextualization enables security analysts to assess threats with a comprehensive picture of the threat landscape[52, 75].
    • Detection–Identifying cyber threats using available indicators, such as CVE [18] and CWE [16]. Detection assists the recognition of both known and emerging threats [114, 32].
    • Correlation–Connecting disparate threats to uncover hidden relationships between attack campaigns and disclose the broader landscape of adversarial community. Correlation enhances situational awareness and strengthens incident investigations by providing a more comprehensive view of threat activities [86, 54, 52].
    • Prioritization– Ranking threats based on severity, exploitability, and potential impact to determine the urgency of remediation. Effective prioritization optimizes resource allocation by focusing on the most critical threats [89, 43].
    • Remediation–Developing and recommending actionable countermeasures, including security patches, mitigation strategies, and responsible advisories. Remediation ensures that organizations can effectively neutralize threats and reinforce their security posture[43, 74].
Within each CTI task, we define distinct analytical targets. For instance, prioritization involves assessing both severity, measured by the Common Vulnerability Scoring System (CVSS) [60], and exploitation potential, measured by the Exploit Prediction Scoring System (EPSS) [29]. The detailed analytical targets for each CTI task are presented in Table 1 and are evaluated extensively in section 5.
2.2 LLM-as-Copilot
Figure 2: Development pipeline of LLM copilots.
Due to their vast inherent knowledge and ability to handle large-scale downstream tasks, LLMs have been extensively adapted for domain-specific applications, an effort referred to as “LLM-as-Copilot” [51, 34, 90]. Unlike the broader concept of an "agent," which is designed to perform human-like behaviors, a copilot specifically emphasizes its role as a human assistant [85, 55, 88]. This distinction is particularly critical in high-stakes domains such as cybersecurity [112, 28], healthcare [4, 98, 108], and finance [107, 111, 70], where expert guidance and decision support are essential.
Training Methodology. As illustrated in Figure 2, the development of LLM copilots follows a pre-training and fine-tuning trajectory. Pre-training leverages large-scale corpora to embed extensive domain expertise into base LLMs through a causal language modeling objective [9, 25, 45, 27]. Subsequently, fine-tuning incorporates environment-aware instructions and utilizes supervised learning [61, 78] to refine the model’s reasoning capabilities for practical applications.
This pre-training followed by fine-tuning approach has proven particularly effective in agent development requiring broad domain expertise, such as existing efforts in AI-assisted programming [2, 91]. Similarly, the interconnected nature of cyber threats necessitates a deep and expansive CTI knowledge. Thus, this work adopts the same development pipeline to enhance LLM capabilities, advancing them toward the creation of an effective and practical cybersecurity copilot.
Prompting. Besides training methodologies, inference-time prompting plays a crucial role in regulating an LLM’s behavior during execution [63, 36]. While developing LLM copilots cannot rely solely on prompting, it provides flexibility by enabling them to interact with real-world resources, such as functional tools [87] or retrieving web-based knowledge [53].
In this work, we enhance CyLens’s inference by implementing 6 NLP modules, integrated through either additional functions or strategic prompting techniques. These enhancements improve CyLens’s reasoning, contextual awareness, and overall effectiveness in CTI analysis.
3 Data Management and Pre-Training
In this section, we describe the collection and preprocessing of threat reports used in pre-training (section 3.1) and introduce a curriculum-based pre-training approach that integrates large-scale cyber threat knowledge into CyLens (section 3.2). The subsequent section (section 4) details the fine-tuning and inference design.
3.1 Threat Data: Collection and Preprocessing
Raw Data Collection. Given that threat data is predominantly represented by CVEs [100, 102], we focus on collecting CVE-centered metadata from several vulnerability databases: MITRE-CVE [14], NVD [59], Exploit-DB [62], CWE [16], CAPEC [17], and ATT&CK [18], including (but not limited to) associated threat actors, TTPs, affected systems, attack infrastructures, and strategical mitigation approaches.
To enrich the collected cyber threat knowledge, we further extract reports and evidence from third-party references such as GitHub commits, Apache [5], and IBM X-Force [39], incorporating diverse cybersecurity activities and countermeasure strategies. These additional sources enhance the comprehensiveness of our dataset by providing real-world exploit details and remediation insights. Table 8 outlines the data sources and the corresponding metadata collected.
Pre-Training Corpus. Next, we preprocess the collected metadata to construct CVE-centric threat reports, which are used as pre-training corpus. As outlined in Algorithm 1, we employ a combination of three approaches to ensure the generation of high-quality reports and diverse layouts:
    • In-Context Learning: We extract sample layouts and phrasing styles from threat reports published by Microsoft [13], Cisco [41], CISA [20], Europol [26], and ENISA [30]. These examples are used as demonstrations $\mathcal{S}$ to guide the generation. For each generation, we select an example report $s\in\mathcal{S}$ and construct a 1-shot prompt, designed to replicate the specific style and layout of $s$.
    • Multi-Agent Generation: For each report generation, we randomly select one LLM from a model set $\Theta$, including GPT-4o, GPT-o1, Gemini-Pro, Pixtral-Large, DeepSeek-R1, or Llama-3.1-405B, to generate threat reports using the collected metadata and a 1-shot prompt. Subsequently, a different LLM is randomly chosen from $\Theta$ to revise the generated reports, introducing variability and enhancing quality by reducing single-model bias.
    • Output Sampling: We adjust two decoding hyperparameters—temperature and top-p—which introduce controlled variability in the outputs. This approach, as studied in [11, 72, 115], helps balance randomness and coherence, ensuring diverse yet meaningful generations.
Input: $\Theta$ – a set of LLMs; $\mathcal{M}$ – collected metadata; $\mathcal{S}$ – example threat reports; $\mathcal{P}_{\texttt{gen}}$ – generation prompt; $\mathcal{P}_{\texttt{rev}}$ – revision prompt;
Output: $D$ – cyber threat corpus;
1 $D\leftarrow\emptyset$ ;
2 foreach metadata for a CVE: $m\in\mathcal{M}$ do
       // Prompt Construction
3       1-shot demo: $s\leftarrow\texttt{RandSelectFrom}(S)$ ;
4       Input prompt: $\mathcal{P}_{m}\leftarrow\mathcal{P}\oplus s\oplus m$;
       // Report Generation and Revision
5       $\theta_{1},\theta_{2}\leftarrow\texttt{RandSelectFrom}(\Theta)$ ;
6       Randomly set temperature $\mathcal{T}_{1},\mathcal{T}_{2}$ and top-p $p_{1},p_{2}$;
7
8      Generation: $d\leftarrow\texttt{OutSampling}(\theta_{1},\mathcal{P}_{m},\mathcal{T}_{1},p_{1})$ ;
9       Input prompt: $\mathcal{P}_{m}\leftarrow\mathcal{P}_{\texttt{rev}}\oplus d$ ;
10       Revision: $d\leftarrow\texttt{OutSampling}(\theta_{2},\mathcal{P}_{m},\mathcal{T}_{2},p_{2})$ ;
11
12      $D\leftarrow D\cup\{d\}$;
13
14 end foreach
15return $D$;
Algorithm 1 Corpus Generation
Due to space constraints, we defer the example reports, input prompts, and the corpus scale to Appendix A.2.
Table 1: Details of the fine-tuning dataset. Each task targets a specific aspect of CTI analysis, incorporating varying dependencies from prior tasks, and involving different NLP modules: TOM–Topic Modeling; NER–Named Entity Recognition; REL–Relation Extraction; RAG–Retrieval-Augmented Generation (using web resources); REA–(any type of) Reasoning; SUM–Summarization.
CTI Task Analytical Target Dependency NLP Module
TOM NER REL RAG REA SUM
❶ Attribution Threat Actor, TTPs, Exploitation Path Initial Evidence
❷ Contextualization Affected System, Attack Infra, Impact Initial Evidence
❸ Detection CVE, CWE Identifier Initial Evidence+❶❷
❹ Correlation Related CVE Identifier Initial Evidence+❶❷❸
❺ Prioritization Severity (CVSS), Exploitation (EPSS) Initial Evidence+❶❷❸
❻ Remediation Patch Tool, Code, Mitigation Advisory Initial Evidence+❶❷❸❹❺
3.2 Curriculum Pre-Training
Next, we pre-train the base LLMs (detailed in Table 9) to embed large-scale cybersecurity knowledge. We design a “curriculum” strategy that progressively introduces the corpus throughout the training process.
Curriculum Design. The “curriculum” refers to the ordered training corpus, structured by the time of threat report publication and the number of tokens in each report. Formally, let $D=\{d_{i}\}_{i=1}^{n}$ represent the corpus, where $d_{i}$ denotes individual threat reports. We first arrange $D$ based on the year and month each CVE was reported. Within each year-month grouping, the reports $\{d_{i}\}$ are further sorted by token length, creating a sequence that progresses chronologically from older to newer reports and by complexity from shorter to longer documents. For simplicity, we continue to denote the ordered corpus as $D$.
Training with Pacing. To enhance convergence and stabilize training, we introduce a pacing function Pace(,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) that governs the portion of the training corpus $D$ utilized during the $t$-th epoch. Formally:
$\textsc{Pace}(D,t)=\begin{cases}D[:|D|\cdot\frac{t}{T_{1}}]&t\leq T_{1}\\ D&T_{1}<t\leq T_{2}\\ D\cup D[\beta\cdot|D|\cdot\frac{t-T_{2}}{T-T_{2}}:]&t>T_{2}\end{cases}$ (1)
The pacing function in Eq. 1 has three stages: (a) Linear Start – In the early epochs ($t\leq T_{1}$), a progressively increasing portion of $D$ is selected for training. (b) Plateau – In the mid stage ($T_{1}<t\leq T_{2}$), as the LLM $\theta$ acquires foundational cyber threat knowledge, the entire corpus $D$ is typically used for training. (c) Reinforced End – In the final epochs ($t>T_{2}$), the corpus $D$ is augmented with the most recent threat data to reinforce training of emerging cyber threats. This augmentation is balanced by a factor $\beta$. Table 9 shows values of $T$, $T_{1}$, and $T_{2}$.
Pre-Training Objective. Identical to the pre-training objective of Llama [92], we employ Causal Language Modeling (CLM), an autoregressive training objective where the LLM predicts the next token in a sequence based on preceding tokens. Let $x_{j}$ denote a token in a threat report $d$ with $L_{d}$ tokens, we have $d=\{x_{j}\}_{j=1}^{L_{d}}$. The CLM objective is formulated as:
$\ell_{\texttt{CLM}}(\theta,D)=-\mathbb{E}_{d\sim\mathcal{D}}\left[\sum_{j=1}^{L_{d}}\log p_{\theta}(x_{j}|x_{<j})\right]$ (2)
Here, the model $\theta$ learns to maximize the likelihood of the next token $x_{j}$ given the preceding context $x_{<j}$, across all tokens in a threat report $d$. This objective enables the model to effectively internalize and represent the knowledge included in the cyber threat corpus $D$.
4 Fine-Tuning, Inference, and Customization
This section details the development of CyLens after pre-training, focusing on task-oriented dataset construction for fine-tuning (section 4.1), modularized inference (section 4.2), and organizational customization (section 4.3).
4.1 Fine-Tuning with Cascading Reasoning
Data Construction. CTI analysis follows a cascading process, which starts with an initial set of evidence and progressively gathers additional information to refine the analysis [64, 81]. To equip LLMs with cascading analytical capabilities, we design the fine-tuning dataset to reflect progressive CTI activities: $D=\{E_{0}\cup(Q_{i},A_{i})_{i=1}^{n}\}$ where $E_{0}$ represents the initial evidence, and each query-answer pair $(Q_{i},A_{i})$ corresponds to a step in the cascading reasoning process:
$(A_{1}|E_{0},Q_{1})\rightarrow(A_{2}|E_{0},A_{1},Q_{2})\rightarrow...\rightarrow(A_{n}|E_{0},A_{1},...,A_{n-1},Q_{n})$
Example 1
Below is a cascading reasoning example.
[…]
Figure 3: An illustration of modularization in CyLens, using the same threat report as described in Example 1.
Table 1 presents the analytical targets for CTI tasks ❶-❻, where each task builds upon the preceding ones. As illustrated in Example 1, identifying CVEs requires first pinpointing threat actors, which subsequently guides the prioritization.
To construct the fine-tuning dataset, we aggregate references from MITRE-CVE, NVD, and ATT&CK. For each reference, we extract available evidence as listed in Table 1 and utilize descriptive reports as the initial evidence set $E_{0}$. We then generate $(Q_{i},A_{i})$ pairs following the format exemplified in Example 1. It is important to note that not all constructed data samples encompass the full pipeline of tasks ❶-❻. Detailed dataset statistics are provided in Appendix B.
Fine-Tuning. To enable CyLens with cascading reasoning capabilities, we fine-tune it using the dataset $D$ described above. The fine-tuning process follows a standard instruction-tuning objective [106, 44], where the model is trained to generate a sequence of answers conditioned on the preceding context. Let $\theta$ represent the parameters of CyLens. The objective function is defined as:
$\begin{split}\mathcal{L}(\theta)=-\mathbb{E}_{(E_{0},Q_{1\sim n},A_{1\sim n})\in D}\ell(E_{0},Q_{1\sim n},A_{1\sim n},\theta)\\ \text{s.t.}\quad\ell=\sum_{i=1}^{n}\log P_{\theta}(A_{i}|E_{0},A_{1},\dots,A_{i-1},Q_{i})\end{split}$ (3)
where $P_{\theta}(A_{i}|E_{0},A_{1},\dots,A_{i-1},Q_{i})$ represents the probability assigned by the model to the correct response $A_{i}$ given the accumulated context up to step $i$.
Table 2: Evaluation results for ❶ Attribution tasks, where "Hist."–historical threats, "0-Day"–zero-day threats. All metrics are scaled to 0-100 (%) for consistency. The boldface and underline highlight the best and second-best results, respectively.
Threat Actor TTP Exploitation Path
Model Accuracy BERT Score Precision Recall F1 Score IoU Accuracy BERT Score
Hist. 0-Day Hist. 0-Day Hist. 0-Day Hist. 0-Day Hist. 0-Day Hist. 0-Day Hist. 0-Day Hist. 0-Day
SEvenLLM 65.94 61.17 81.25 79.83 73.44 68.62 76.85 71.53 75.11 70.04 60.14 53.90 73.56 78.42 82.19 80.65
Lily-Cyber 63.62 54.58 78.25 75.27 64.71 55.32 70.17 64.62 59.65 67.28 42.50 50.69 76.25 70.38 82.16 82.43
ChatGPT-4o 48.33 40.31 76.02 73.63 74.46 77.42 81.53 80.52 77.83 78.94 63.71 65.21 67.76 69.74 76.15 79.90
ChatGPT-o1 57.36 55.34 77.93 78.69 81.26 72.49 83.18 82.90 82.21 77.35 69.79 63.06 78.82 72.78 77.22 78.06
Gemini-Pro 61.06 57.43 81.24 78.63 78.11 79.36 81.19 82.24 79.62 80.77 66.14 67.75 62.82 69.21 75.84 74.43
Llama-405B 34.85 24.62 74.20 71.35 72.46 60.29 78.93 71.08 75.56 65.24 60.72 48.41 41.64 25.72 74.54 71.58
DeepSeek-R1 57.85 55.26 80.39 79.92 73.24 71.51 76.43 79.85 74.80 75.45 59.75 60.58 65.71 63.04 78.27 77.91
Claude-Opus 30.12 22.64 73.51 72.96 69.82 54.84 77.95 75.26 73.66 63.45 58.30 46.46 58.91 52.64 76.46 72.83
CyLens-1B 82.81 79.52 83.84 81.72 86.15 83.22 85.72 87.44 85.93 85.28 75.34 74.33 85.28 80.61 84.62 81.55
CyLens-8B 87.58 83.74 86.72 85.39 90.67 88.70 93.75 92.74 92.18 90.68 85.50 82.94 95.71 87.79 88.45 85.21
CyLens-70B 96.53 88.29 87.93 84.52 89.63 87.57 88.26 85.46 88.94 86.50 80.08 76.21 91.53 86.06 87.32 87.03
4.2 Modularized Inference
At each reasoning step $(E_{0},A_{1},\dots,A_{i-1},Q_{i})\rightarrow A_{i}$ during inference, we employ various NLP techniques to modularize the analytical process. Table 1 summarizes the NLP techniques used for each task. These techniques are implemented through the following approaches:
Topic Modeling. We implement topic modeling to interpret users’ intentions regarding the tasks they aim to perform. Specifically, we use BERTopic, a transformer-based package that eliminates the need for manually setting hyperparameters, as required in conventional methods [7, 50]. Given a threat report $E_{0}$, we implement a two-step approach: first, based on CTI tasks ❶-❻, and then, based on specific analytical targets, as outlined in Table 1.
Example 2
Topic modeling (Figure 3-(b)).
Using the same example as in 1, the topic modeling step first determines the task as threat attribution based on $Q_{1}$. It then further specifies the analytical target as the threat actor within the threat attribution task.
Named Entity Recognition (NER) and Relation Extraction. Next, we extract entities relevant to the identified topics while simultaneously identifying relationships among them. We employ the multi-agent generation approach described in section 3.1, utilizing few-shot prompting, where task-specific entities and their relationships are included as demonstrations within the prompt. This approach facilitates a structured understanding of cybersecurity threats, enhancing the accuracy and contextual relevance of extracted threat intelligence.
Retrieval Augmented Generation (RAG). We implement RAG to retrieve relevant web documents through web screening to enrich threat reports. This is especially useful for zero-
Example 3
NER and relation extraction (Figure 3-(c)).
Using Example 1, the NER identifies key entities "APT28" and "C2 communication" within the threat report. Simultaneously, relation extraction determines the "is associated with" relationship between these entities.
day threat hunting, where web-scale CTI intelligence—such as reported threat groups, affected products, and potential patches—is continuously updated. Then, we prompt GPT-o1 to rank the retrieved documents based on their relevance to the threat report $E_{0}$. To control latency, we implement a cache to store retrieved documents. Caching documents are indexed by topic modeling results (i.e., CTI tasks and analytical targets) and specific entities (e.g., APT28 or C2 communication).
Reasoning and Summarization. We implement reasoning and summarization through prompting in CyLens, where reasoning is applied when deeper analysis or multi-step thinking is required. For example, reasoning is used to analyze severity level (e.g., medium) based on gathered evidence. Summarization is then applied to synthesize insights, tailoring the output to a specific CTI task by leveraging results from topic modeling, NER, relation extraction, and RAG.
Cascading Reasoning in Inference. Since CTI tasks exhibit dependencies (Table 1), we minimize latency by concurrently executing topic modeling, NER, relation extraction, and RAG across all prerequisite tasks. Reasoning and summarization then proceed sequentially, following task dependencies to generate structured, step-by-step findings and conclusions.
4.3 Organizational Customization
Although CyLens is developed using massive CTI knowledge, it also enables customization to meet specific organizational interests on particular threat categories, specific products, or assisting an entire organization. Such customization is achieved by further fine-tuning LLMs but specifying the scope of data—for example, concentrating on cyber threats related to a single organization. This targeted customization enhances the relevance of CTI in meeting organizational needs.
We demonstrate CyLens’s potential for tailored customization in section 5.6. Through experiments, we highlight its adaptability and flexibility in addressing diverse CTI challenges across different organizational settings.
5 Evaluation Results
This section systematically evaluates CyLens’s performance on CTI tasks ❶-❻ with the following settings:
Dataset. To ensure practicality, we directly assess threats from NVD [59], MITRE [14], and Exploit-DB [62], categorizing into two groups: (1) Historical – threats reported between 2015 and 2024. (2) Zero-day – newly reported threats from the past two months during our evaluation period in 2025.
For each threat, we compile source references from [14, 59] into structured reports and prompt LLMs for analysis. The prompt structure is illustrated in Figure 4. For different tasks, we modify the system prompt and questions accordingly.
To ensure fairness, no test threats were included in training. Statistics and detailed settings are provided in Appendix C.
Baseline. We primarily consider two groups of baselines: (1) State-of-the-art (SOTA) CTI agents, SevenLLM [44] and Lily-Cyber [84], specifically developed for CTI, and (2) Industry-leading general-purpose LLMs, such as ChatGPT and Gemini, which are widely used. Additionally, we investigate the impact of ablation baselines in section 6.1.
Metrics. We use different evaluation metrics depending on the specific CTI task. The definitions, strategies, and rationale for using these metrics are provided in Appendix C.2.
Table 3: Evaluation results for ❷ Contextualization tasks, where "Hist."–historical threats, "0-Day"–zero-day threats. All metrics are scaled to 0-100 (%) for consistency. The boldface and underline highlight the best and second-best results, respectively.
Affected System (w/ Version, if available) Attack Infrastructure Impact
Model Precision Recall F1 Score IoU Precision Recall F1 Score IoU BERT Score
Hist. 0-Day Hist. 0-Day Hist. 0-Day Hist. 0-Day Hist. 0-Day Hist. 0-Day Hist. 0-Day Hist. 0-Day Hist. 0-Day
SEvenLLM 24.76 13.08 18.57 6.53 21.22 8.71 11.87 4.55 43.17 40.28 53.62 45.92 47.83 42.92 31.43 27.32 75.92 77.64
Lily-Cyber 27.73 22.42 19.62 14.57 22.9 8 17.66 12.98 9.69 44.73 35.16 56.82 53.13 50.06 42.32 33.38 26.84 73.02 76.81
ChatGPT-4o 53.76 48.62 41.80 36.09 47.03 41.43 30.75 26.13 54.51 47.78 62.55 57.22 58.25 52.08 41.10 35.20 76.52 78.16
ChatGPT-o1 58.74 55.49 58.61 47.37 58.67 51.11 41.52 34.33 61.72 50.33 65.23 63.26 63.43 56.06 46.44 38.95 78.91 80.24
Gemini-Pro 35.62 33.17 41.52 34.50 38.34 33.82 23.72 20.35 42.35 34.35 31.17 22.62 35.91 27.28 21.88 15.79 74.85 72.62
Llama-405B 11.46 8.12 21.02 10.74 14.83 9.25 8.01 4.85 31.29 28.74 27.08 20.35 29.03 23.83 16.98 13.53 72.24 68.33
DeepSeek-R1 36.71 31.88 54.63 47.62 43.91 38.19 28.13 23.60 48.92 33.96 40.92 31.87 44.56 32.88 28.67 19.68 74.26 72.41
Claude-Opus 31.26 29.82 27.35 22.41 29.17 25.59 17.08 14.67 44.65 41.71 48.90 46.48 46.68 43.97 30.44 28.18 76.32 74.15
CyLens-1B 78.64 74.45 81.24 77.26 79.92 75.83 66.55 61.07 85.73 82.24 87.62 86.61 86.66 84.37 76.47 72.96 83.75 85.26
CyLens-8B 87.91 85.82 91.86 88.12 89.84 86.95 81.56 76.92 87.46 85.71 92.83 88.17 90.07 86.92 81.93 76.87 90.03 88.31
CyLens-70B 94.54 88.17 96.04 93.75 95.28 90.87 90.99 83.28 96.37 90.16 97.69 93.18 97.03 91.65 94.22 84.58 87.33 85.84
5.1 Threat Attribution
We begin by attributing the origin of cyber threats with three targets: threat actor, TTP, and exploitation path, as outlined in Table 1. Our results yield several research findings (RF):
RF1: Specialized Development is Necessary. As shown in Table 2, industry-leading LLMs, despite their broad applicability across diverse tasks, exhibit limited performance in CTI analysis. For instance, ChatGPT-4o has achieved only
SYSTEM You are a cybersecurity threat analyst specializing in cyber threat intelligence (CTI). Your task is to assess potential attack tactics, techniques, and procedures (TTPs), and provide an intelligence assessment following structured industry frameworks such as $\cdots$ USER <Report Title> "Emerging APT Group Exploiting Zero-Day Vulnerabilities in Cloud Infrastructure." <Report Excerpt> Recent intelligence reports indicate that an Advanced Persistent Threat (APT) group has been exploiting an unpatched zero-day vulnerability in cloud service environments. The group is using spear-phishing emails to gain initial access, deploying a custom malware variant named ‘CloudPhantom,’ and leveraging compromised API keys to move laterally $\cdots$ <Question> Identify the relevant TTPs used in this attack. Let’s think step-by-step, first, determine CTI tasks among $\cdots$ ASSISTANT <CTI Task><Analytical Target><Key Entities><Entity Relations><Retrieved Evidence and Source>$\cdots$<Final Answer>
Figure 4: Prompt example for TTP analysis. For both baselines and CyLens, we prompt them to perform step-by-step analysis following the NLP modules defined in Table 1. For tasks with prior dependencies (e.g., remediation), we instruct LLMs to first complete the prerequisite tasks before proceeding.
40.31% accuracy in threat actor attribution, significantly lower than CyLens-1B (79.52%) and CyLens-8B (83.74%) for the same task. This gap underscores the necessity of specialized development, as detailed in section 3 and section 4. Tailoring models specifically for CTI significantly reduces parameter requirements—CyLens-1B and 8B are far more lightweight compared to the 500B–1T parameters of GPT-4o.
RF2: Large model size is not always beneficial. Interestingly, scaling up model size does not always translate to improved CTI effectiveness. For instance, CyLens-8B achieves 93.74% recall in TTP attribution, outperforming CyLens-70B (88.26%), suggesting diminishing returns despite undergoing the same development process. This indicates that larger models do not alaways yield better performance in CTI tasks, emphasizing the importance of domain-specific optimizations over sheer model size.
RF3: Task-specific capability varies across baselines. As shown in Table 2, baseline models exhibit varying levels of effectiveness across different CTI tasks. For instance, GPT-o1 performs obviously worse in attributing threat actors compared to exploitation paths. A deeper case study reveals that baseline LLMs struggle to accurately interpret CTI-specific terminology related to threat actors. For example:
Case Study 1
When analyzing Log4Shell (a known vulnerability) with Mirai (a botnet) as the threat actor, GPT-o1 describes the attacker as an “outside attacker”, failing to precisely identify Mirai. This demonstrates a fundamental gap in attribution accuracy for complex cyber threats.
On the other hand, GPT-o1 achieves $>$78% accuracy in analyzing exploitation paths, as this task primarily involves summarizing existing evidence instead of complex reasoning.
Case Study 2
In cases where specificity is less critical, a rough description—such as "JNDI lookup and remote code fetching"—sufficiently summarizes the Log4Shell exploitation path. While a more precise description would be: "The Log4j library contacts the attacker’s malicious LDAP server, which then responds with a reference to a remote Java class file containing malicious code," the broader summary still effectively conveys the core mechanism.
This indicates that baseline models perform better in tasks with higher generation flexibility but face challenges in tasks requiring precise CTI expertise and structured analysis.
RF4: CyLens Performs consistently across analyses. Unlike baseline models, CyLens maintains consistent performance across different analytical targets, highlighting the importance of incorporating comprehensive expertise during development. This balanced capability ensures that CyLens does not exhibit biased performance or knowledge gaps throughout the CTI lifecycle. The same conclusion is further reinforced by subsequent evaluations.
5.2 Threat Contextualization
We continue evaluating CyLens in contextualizing affected systems, attack infrastructure (e.g., domains and servers), and impact. Besides the similar findings as outlined in RF1-RF4, our evaluations derive complementary research findings:
RF5: Zero-day threat hunting is not always harder. Both Table 2 and 3 reveal that analyzing zero-day threats does not always result in worse performance compared to historical threats. For instance, when assessing the impact of cyber threats, LLMs such as CyLens-1B and SevenLLM sometimes achieve slightly better relevance in their generated analyses. We attribute this to the nature of specific CTI tasks: For impact analysis, the contextualized understanding of previously known threats allows LLMs to provide more insightful assessments, particularly when certain zero-day threats resemble threats they have already encountered. However, for tasks such as identifying affected systems, attack infrastructures, or threat actors, zero-day threats often introduce novel and previously unseen products or malware, making analysis more challenging compared to historical threats.
RF6: Larger LLMs tend to be wordy. Through detailed case study, we observe that larger LLMs tend to generate verbose responses, leading to lower text similarity as measured by BERT score. For example:
Case Study 3
CyLens-8B summarizes an impact analysis concisely: "The March 2022 cyberattack by the Lazarus Group resulted in the theft of \$620 million from the Ronin Network, significantly impacting the cryptocurrency market and…" This closely aligns with the reference [101], maintaining factual accuracy while remaining succinct. In contrast, CyLens-70B generates a more verbose response: "In March 2022, the notorious Lazarus Group orchestrated a sophisticated cyber heist, infiltrating the Ronin Network’s defenses to abscond with a staggering \$620 million in cryptocurrency. This audacious act not only sent shockwaves through the digital currency markets, causing a temporary plunge in asset values, but also undermined investor confidence…" The latter output is overly wordy and includes sentiment-heavy language (e.g., "notorious," "sophisticated," "staggering," "audacious"), making it less aligned with the original reference.
This finding aligns with RF2, reinforcing that larger models are not always more effective for CTI tasks. However, to balance overall performance across all CTI tasks, a systematic evaluation is necessary to determine the optimal model scale.
RF7: CTI agents should align with real-world practice. Despite the strong performance of SOTA agents (SevenLLM and Lily-Cyber) on certain tasks (e.g., analyzing TTPs), they struggle with others, such as attack infrastructure analysis. We attribute this to a misalignment between their training focus and CTI applications. For instance, SevenLLM is trained with a broad emphasis on NLP tasks (e.g., NER), while Lily-Cyber is developed using hand-crafted cybersecurity and hacking-related data pairs [84]. However, these training paradigms often conflate TTPs and threat actors, rather than treating them as distinct analytical components. This lack of granularity can hinder precise CTI reasoning, particularly for tasks that require structured analysis of attack properties.
Although CTI lacks a universally standardized taxonomy, structuring tasks in alignment with real-world CTI databases (e.g., MITRE-CVE and NVD) provides a more practical and reliable foundation for cyber threat analysis. This highlights the necessity of LLMs specifically designed for CTI, rather than relying on models with indirect exposure to cybersecurity knowledge or misaligned customization that does not fully reflect real-world CTI practices.
Table 4: Evaluating ❺ Prioritization performance using CVSS 3.x metrics [60]: AV–Attack Vector, AC–Attack Complexity, PR–Privileges Required, UI–User Interaction, S–Scope, C–Confidentiality Impact, I–Integrity Impact, A–Availability Impact, Base–CVSS Base Severity. We report the accuracy of correctly categorizing each metric (e.g., Attack Vector: Network). The boldface and underline highlight the best and second-best results, respectively.
Model Historical Threat Zero-Day Threat
AV AC PR UI S C I A Base AV AC PR UI S C I A Base
SL 22.51 82.64 35.04 54.21 56.87 30.14 21.62 54.67 32.75 14.76 76.85 24.46 32.27 46.83 33.70 35.62 51.84 27.58
LY 54.75 65.83 22.12 76.91 53.56 43.17 37.85 42.73 43.56 47.52 64.31 19.75 59.17 58.75 37.28 33.54 45.19 36.61
4O 76.85 92.91 61.84 80.78 66.51 27.05 20.43 34.21 58.96 77.81 84.55 56.78 62.35 68.14 22.56 31.05 28.49 46.51
O1 82.51 93.36 70.57 88.21 72.53 51.69 24.22 46.73 69.25 86.83 92.74 63.17 77.14 76.83 45.30 47.26 54.03 63.71
GM 82.07 90.71 46.42 73.72 61.24 44.75 39.80 38.19 52.51 67.38 83.26 31.07 56.34 55.62 31.27 36.56 42.93 39.81
LA 37.18 56.25 67.24 44.21 48.67 22.13 18.69 29.32 24.36 12.61 57.63 54.53 39.81 28.75 10.33 17.72 11.36 17.45
DS 64.52 87.21 54.52 87.36 69.54 46.27 44.18 40.21 62.64 58.82 82.01 46.42 82.86 51.35 35.26 40.96 38.74 53.26
CD 53.76 66.25 63.87 56.32 62.71 35.06 41.28 37.93 49.07 52.16 55.86 66.46 49.39 57.24 30.04 22.75 28.39 36.88
C1 86.26 93.32 88.27 94.55 95.51 81.10 79.63 84.23 88.17 84.71 90.83 86.06 96.64 90.32 76.64 73.59 81.62 83.44
C8 94.74 98.85 95.53 99.73 97.63 93.39 90.33 95.49 96.86 93.27 99.31 94.42 97.65 94.52 85.43 88.47 86.54 91.57
C70 94.33 100.0 97.56 97.46 98.15 94.74 92.26 95.83 97.57 92.69 100.0 97.71 95.48 93.26 90.97 94.53 92.66 93.16
5.3 Detection and Correlation
We now evaluate CyLens and baselines on ❸ detection and ❹ correlation tasks. In detection, the objective is to evaluate how effectively LLMs can identify relevant threat identifiers (CVEs and CWEs) from a given threat report. This step is crucial for enabling subsequent CTI activities, such as threat correlation and remediation. In correlation, we extend the evaluation by examining how LLMs associate other threats (CVEs) using different strategies: (i) Direct correlation–directly correlate additional CVEs. (ii) CVEs as intermediates–first detect CVEs from threat reports (from ❸) and then use them for correlation. (iii) CWEs as intermediates–first identifying CWEs (i.e., weakness categories, from ❸) and using them for correlations. The latter two strategies leverage intermediate evidence (CVEs or CWEs identified in the detection step) to uncover the threat network.
Our evaluations yield the following new research findings:
Figure 5: Evaluation results on ❸ Detection tasks using historical threats. The presentation order of LLM names (top-down) follows the same sequence as in Tables 2 (top-down).
RF8: CyLens demonstrates strong detection relevance. Figure 5-(a)(b) shows the hit ratio for the top 10 detected threat identifiers (CVEs, CWEs). Identifying CVEs is more challenging due to their vast number ($>$200k) compared to CWEs ($\sim$900). However, we also observe that CyLens  is able to detect more relevant CVEs such as those that belong to the same category or affect the same products. For example:
Case Study 4
Given the threat report: "…Apache HTTP Server allows remote attackers to execute arbitrary code. The issue arises from improper handling of user-supplied input in HTTP request headers…" corresponding to CVE-2021-42013. LLMs present different detections: GPT-o1 and DeepSeek-R1 incorrectly output CVE-2020-0609 and CVE-2022-2296, both of which are unrelated to Remote Code Execution or Apache. CyLens-70B gives CVE-2023-25690, a highly relevant HTTP Request Smuggling attack on the Apache Server.
From the above case study, although CyLens-70B does not match the exact CVE identifier, it effectively aligns with the evidence provided in the threat report and demonstrate a strong threat detection capability. This effectiveness is further supported by the results shown in Figure 5-(c)(d).
RF9: CyLens demonstrates robust threat correlation. Table 5 presents the hit ratio for the top 10 correlated threats (CVEs) as reported in MITRE-CVE [14]. We evaluate three correlation strategies: (i) Direct correlation, (ii) CVEs as intermediates, and (iii) CWEs as intermediates, as described above. We have corresponding observations:
Table 5: Evaluation results (HIT@10) on ❹ Correlation task using historical threats. The order of LLM names (left-to-right) follows the same sequence as in Tables 2 (top-down).
SL LY 4O O1 GM LA DS CD C1 C8 C70
Direct 0.6 0.0 3.5 8.1 2.2 0.0 0.0 0.4 52.7 74.3 77.5
w/ CVE 7.3 4.5 11.6 25.4 6.7 0.0 4.3 12.5 60.4 79.0 81.2
w/ CWE 0.0 0.0 0.0 0.8 0.0 0.0 0.0 0.0 41.7 70.8 71.6
    • (i)
      While baseline models largely struggle to correlate CVEs, CyLens demonstrates significant effectiveness (e.g., 77.5% Hit@10 with CyLens-70B).
    • (ii)
      Incorporating CVEs as intermediate evidence improves correlation performance. However, baseline models tend to propagate erroneous CVEs from the prior detection step, which misguides subsequent threat correlations.
    • (iii)
      Incorporating CWEs degrades correlation performance due to their inherent ambiguity. Since a CWE can broadly relate to many CVEs, it introduces ambiguity, impeding accurate threat correlation. Despite this challenge, CyLens remains robust (e.g., 71.6% Hit@10 with CyLens-70B) in incorporating ambiguous intermediate evidence while maintaining strong correlation effectiveness.
5.4 Threat Prioritization
Real-world cyber threats often co-occur across cyber events [77, 73], necessitating that CTI teams distinguish threat severity and prioritize subsequent patches and mitigations. Our evaluation focuses on two prioritization measurements: severity level and analysis of exploitation dynamics.
Severity Level. Threat severity can be assessed from threat reports even when CVE identifiers are not disclosed or when zero-day threats have not yet been assigned CVEs. We evaluate severity using 8 metrics defined in CVSS 3.x [60], which collectively capture a threat’s exploitability and impact. Each severity metric is treated as a classification task. For example, the Attack Vector metric has four predefined categories: Network, Adjacent Network, Local, and Physical. In cases where a threat report lacks sufficient evidence for CVSS metrics, we introduce an additional category, "N/A", for each metric.
Furthermore, we classify the base severity level from four categories: Low, Medium, High, and Critical. This classification aggregates the ratings from 8 CVSS metrics, providing an overall severity assessment aligned with CVSS standards.
RF10: CyLens maintains stable effectiveness. Table 4 shows the accuracy of all models. While baselines vary widely across CVSS metrics (e.g., GPT-o1 scores 92.91% on Attack Complexity but only 20.43% on Integrity Impact), CyLens remains consistently strong. Notably, CyLens-8B achieves $>$90% accuracy across 8 metrics for historical threats, demonstrating its reliability in handling different severity factors.
This stability comes from CyLens ’s ability to adapt to different evidence types. Each CVSS metric relies on specific context—Privileges Required (PR) depends on access control policies, while Attack Complexity (AC) is influenced by real-world constraints and attacker skill level. By effectively integrating diverse evidence sources, CyLens reduces inconsistencies in severity analyses, which aligns with RF4.
Additionally, we find that zero-day threats are not always harder to analyze, aligning with RF5. Some evidence, like some user actions in exploitations, is already known from past threats, facilitating assessment on newly emerging threats.
We also observe that larger LLMs do not always perform better, consistent with RF6. While bigger models tend to provide more detailed reasoning, this can sometimes lead to inaccurate analysis. For example, CyLens-70B does not always outperform CyLens-8B across all severity metrics.
Exploitation Dynamics. Besides severity analysis using CVSS, we also evaluate CyLens’s effectiveness in analyzing dynamic exploitation behaviors. Specifically, we use the EPSS score [29] and conduct two types of analysis: (1) pinpointing historical exploitation at an earlier timestamp and (2) forecasting future exploitation at a later timestamp.
Figure 6: With $\mathcal{T}_{0}$ as the “current” timestamp (varying among threats), we compute RMSE for historical EPSS scores over the past three months (interpolation) and future EPSS scores for the next three months (prediction) w.r.t. $\mathcal{T}_{0}$.
For test threats with more than six months of EPSS records, we set a reference timestamp $\mathcal{T}_{0}$ for each cyber threat (e.g., 01/01/2024) and collect EPSS reports from $\mathcal{T}_{0}$-3 months (e.g., 10/01/2023) to $\mathcal{T}_{0}$+3 months (e.g., 04/01/2024). However, EPSS reports between $\mathcal{T}_{0}$-3 to $\mathcal{T}_{0}$ are included in the inference prompt as prior knowledge, similar to the prompt structure illustrated in Figure 4.
We compare CyLens-1B/8B against four baseline models and compute RMSE between the interpolated (historical) or predicted (future) EPSS scores and the actual scores (scaled from 0% to 100%). To ensure accurate evaluation, we match each EPSS score to the closest available recorded score. For example, if an EPSS score is not reported exactly on 10/01/2023, we use the closest available date (e.g., 09/25/2023) from EPSS reports [14] as the ground truth.
RF11: LLMs vary in interpolation and prediction. As shown in Figure 6, LLMs perform slightly better in interpolating historical EPSS scores (lower RMSE) than in predicting future scores. However, we also observe that local LLMs, such as SevenLLM, perform obviously worse in both interpolation and prediction due to their static inherent knowledge, which lacks adaptability to evolving cyber threats. On the contrary, CyLens, despites its local deployment, performs similar to API-based LLMs (e.g., Gemini), which demonstrates the necessity of equipping LLMs with domain expertise. This enables LLMs to adapt to the dynamic nature of CTI tasks.
RF12: Contextual knowledge is essential. Beyond Figure 6, we observe that some cyber threats undergo dramatic shifts in EPSS scores due to emerging exploitation activities.
Case Study 5
For instance, CVE-2024-27956 saw a sharp increase in EPSS on 02/15/2025, jumping from 0.31% to 55.37% [19]. No LLM was able to accurately predict this trend without contextual knowledge of real-world events.
This implies the importance of not only fitting to EPSS trends but also incorporating real-world evidence in real time to enhance CTI adaptability.
5.5 Remediation
Table 6: Evaluation results on ❻ Remediation. We evaluate whether LLMs can correctly generate: (i) patching or mitigation tools – measured by accuracy, (ii) code patches – measured by syntactic similarity, (iii) patching or mitigation methods – measured by BERT Score, and (iv) reference advisories – measured by accuracy.
Model Tool Use Code Patch Methodology Advisory
Hist. 0-Day Hist. 0-Day Hist. 0-Day Hist. 0-Day
SL 87.81 79.65 81.83 76.62 82.95 82.48 56.74 43.25
LY 79.63 57.17 75.96 73.60 80.75 78.53 48.71 46.63
4O 86.62 74.51 82.70 81.05 84.42 82.51 76.83 71.75
O1 90.08 86.65 84.47 82.86 85.53 85.25 81.26 77.79
GM 76.83 53.22 83.14 82.47 84.79 80.32 66.72 43.25
LA 75.71 69.20 81.23 78.46 83.26 82.59 49.53 38.17
DS 73.70 70.21 79.27 77.51 83.64 82.55 65.43 47.92
CD 77.54 73.48 76.19 73.64 82.27 81.92 54.85 47.35
C1 92.03 88.84 85.42 83.71 87.83 84.52 88.74 82.25
C8 93.83 92.74 87.48 87.36 89.16 88.37 92.32 88.54
C70 97.62 97.23 90.66 86.39 91.20 88.04 95.34 92.26
Finally, we evaluate CyLens’s capability in generating accurate remediation strategies, focusing on four aspects: suggesting the correct tool for patching or mitigation (e.g., OpenSSL for encryption), generating available code patches (e.g., GitHub commits), proposing appropriate mitigation methodologies (e.g., hardening an OS), and identifying the correct advisory to consult (e.g., NVIDIA Security Advisory).
RF13: CyLens is effective across all remediation tasks. Consistent with RF4 and RF10, we demonstrate CyLens’s effectiveness in handling various remediation objectives as shown in Table 6, including both identification-based (suggesting tools and advisories) and generation-based (producing code patches or mitigation strategies) approaches. In contrast, baseline models, lacking CTI-specific expertise, often exhibit imbalanced performance across remediation tasks. For example, Gemini performs well in generating code patches but struggles to identify the correct advisories.
Figure 7: Evaluation results across multiple CTI tasks based on different levels of organizational interests: (a)-(d) product-level, (e)-(h) vendor-level, and (i)-(l) threat category-level. The evaluation metrics used are accuracy for threat actor attribution, F1 score for attack infrastructure contextualization, accuracy for CVSS base severity classification, and accuracy for third-party advisory suggestion. Model names are SL: SevenLLM, O1: ChatGPT-O1, GM: Gemini-Pro, C1/C8: CyLens-1B/8B.
RF14: Large-scale expertise is crucial. Effective remediation often relies heavily on existing cybersecurity practices [43]. Therefore, embedding large-scale historical remediation knowledge is crucial when developing LLMs for CTI applications. Even for zero-day threats, well-trained LLMs can leverage their inherent knowledge to derive useful insights, improving response effectiveness even in cases where no explicit remediation has been documented.
5.6 Organizational Customization
Previous experiments demonstrate CyLens’s capability in analyzing a wide range of cyber threats, providing a strong foundation for addressing specific organizational needs. Given that threats of particular interest—such as those associated with specific products, vendors, or threat categories—may lack sufficient volume for fine-tuning [58, 35], it is essential to first develop LLMs with comprehensive CTI analysis capabilities. These LLMs can then be incrementally customized to align with specific organizational needs.
We categorize organizational needs into three levels:
    • Product-Level – Focusing on cyber threats affecting specific products, a common approach among security teams in large enterprises managing product-specific CTI [21].
    • Vendor-Level – Targeting threats that impact an entire vendor (i.e., company), a perspective often adopted by small and medium-sized businesses or industry groups [67].
    • Category-Level –Concentrating on specific threat categories, which is particularly relevant to research institutions conducting specialized cybersecurity studies [49].
Settings. To cover diverse organizational needs, we collect threats associated with (1) specific products (Acrobat, Chrome, or Fedora), (2) specific vendors (openSUSE, GitLab, or Dell), and (3) specific threat categories (Overflow, DoS, or XSS). Thus, we collect a total of nine groups of threats, with each group comprising more than 1,000 threat reports. We allocate 50% of these threat reports—sourced from third-party references in the MITRE-CVE [15], NVD, and Exploit-DB [62] websites [59]—for independently fine-tuning CyLens-1B/8B, while reserving the remaining for evaluation. We compare CyLens to the SOTA model, SevenLLM, and two industry-leading LLMs, GPT-o1 and Gemini. Figure 7 presents the results across multiple CTI tasks.
RF15: CyLens is adaptable to diverse organizational needs. From Figure 7, we observe that CyLens maintains high effectiveness even with independent customization (fine-tuning) across different groups of cyber threats. For instance, CyLens-8B achieves nearly 100% accuracy in providing remediation advisories for various vendors and threat categories. This result further demonstrates its adaptability despite the diverse nature of threats across products, vendors, and categories—factors that significantly impact baseline LLMs, causing their performance to fluctuate. The finding implies CyLens ’s superior capability to generalize while maintaining high effectiveness in tailored needs.
6 Discussion
6.1 Ablation Study
We conduct ablation studies to assess the impact of different components within CyLens. Table 7 presents the results, where we independently exclude pre-training (PT), fine-tuning (FT), and four optional modules: topic modeling (TOM), named entity recognition (NER), relation extraction (REL), and retrieval-augmented generation (RAG). Notably, reasoning (REA) and summarization (SUM) cannot be excluded, as they are essential steps in text generation.
Table 7: Ablation study on CyLens-8B by selecting four representative CTI tasks: (1) attributing threat actors (measured by accuracy), (2) contextualizing attack infrastructure (measured by F1 score), (3) classifying CVSS base severity levels (measured by accuracy), and (4) suggesting advisories for remediation (measured by accuracy).
Ablation Threat Actor Attack Infra CVSS Base Advisory
Hist. 0-Day Hist. 0-Day Hist. 0-Day Hist. 0-Day
w/o PT 66.22 62.31 70.05 61.44 53.62 45.90 57.25 54.36
w/o FT 82.42 70.15 86.47 78.31 86.48 61.75 82.56 77.23
w/o TOM 84.23 81.17 88.46 82.24 95.12 88.82 87.23 85.44
w/o NER 71.12 70.19 82.83 72.63 94.52 91.42 82.53 80.65
w/o REL 87.21 83.09 88.18 86.21 95.67 91.25 83.12 82.72
w/o RAG 84.17 80.26 86.53 82.71 95.23 89.64 85.14 83.42
We observe that pre-training has the most significant impact, as it embeds massive CTI knowledgeinto the model. While fine-tuning is more lightweight, its influence remains non-negligible in refining CyLens ’s CTI capabilities. Additionally, the impact of different modules varies across tasks. For instance, classifying CVSS severity levels does not heavily depend on NER and REL, so removing these modules has minimal effect. However, tasks such as threat actor attribution or advisory suggestion rely on accurately identifying key objects (e.g., malware or third-party vendors), making NER and REL critical for performance. These findings reinforce the necessity of incorporating all proposed techniques to achieve an effective and well-rounded system design for CTI applications.
6.2 Adapting to Computational Constraints
Similar to Meta’s Llama series, which offers models at different scales [56, 57, 92], CyLens is available in three sizes: 1B, 8B, and 70B, allowing it to accommodate varying computational constraints. While CyLens-1B is less effective compared to the 8B and 70B versions, its lightweight design is crucial for resource-limited deployments.
In scenarios where computational resources are highly constrained, having a functional LLM is often more practical than striving for state-of-the-art performance. Notably, despite its smaller size, CyLens-1B still outperforms industry-leading LLMs in most CTI tasks, reinforcing its value as an efficient yet capable model for cybersecurity applications.
6.3 Privacy-Preserving Deployment
A key advantage of customizing local LLMs is enhanced privacy preservation. Unlike API-based services such as OpenAI’s GPT series, which inherently risk exposing user intentions and confidential data, local models ensure that sensitive information remains under organizational control. API-based solutions may not only transmit proprietary data to external servers but could also use it for model training and further development, raising concerns about data security and regulatory compliance [65, 66, 80, 99, 113].
By developing localized, task-specific LLMs, organizations gain full control over model behavior, customization, and data security. Such privacy-preserving deployment is particularly critical in high-stakes fields like CTI, where threat intelligence, incident reports, and security strategies must remain confidential to prevent adversarial exploitation and ensure compliance with data protection regulations.
6.4 Beyond Specific Needs
Real-world cyber threats are rarely confined to specific products, vendors, or categories [77]. Instead, they manifest within a complex and interconnected ecosystem where vulnerabilities, exploits, and attack methodologies transcend traditional boundaries. This interdependence underscores the necessity of developing CTI solutions that go beyond addressing specific organizational needs (e.g., tailored threat intelligence) and are also capable of capturing the broader, multifaceted threat landscape.
7 Conclusion
This paper introduced CyLens, a LLM-powered copilot for cyber threat intelligence. CyLens streamlines the entire threat management lifecycle, encompassing tasks from threat attribution to remediation. It assimilates knowledge from a vast number of threat reports and utilizes a retrieval module for real-time information, enabling it to adeptly handle both historical and zero-day threats. Rigorous evaluations have demonstrated that CyLens consistently outperforms industry-leading LLMs and SOTA agents across multiple CTI tasks. Its adaptability shines through its customization capabilities, which can be tailored to specific threat categories or organizational requirements. Additionally, the release of its training datasets fosters further research in the field.
Appendix A Pre-Training: Complementary Information
Table 8: The data sources we utilize and the corresponding evidence metadata collected from each database.
CTI Task Analysis Target Source Database
CVE NVD CWE ATT&CK CAPEC Exploit DB Third-Party Ref.
Threat Actor
TTPs
Exploit Path
Affected System
Attack Infra
Attack Impact
CVE Identifier
CWE Identifier
CVE Identifier
CVSS Metrics
EPSS Records
Tool Use
Code Patch
Methodology
Advisory
Table 9: Base LLM assignment and settings of CyLens’s development, including pre-training and fine-tuning. “Cp” – Corpus; “Tk” – token.
CyLens Base LLM Pre-Training
# Cp # Tk $T$ $T_{1}$ $T_{2}$ $\beta$
CyLens-1B Llama-3.2-1B 28K 0.27B 4 3 3 1.0
CyLens-8B Llama-3.1-8B 115K 1.1B 10 5 5 1.0
CyLens-70B Llama-3.3-70B 271K 2.5B 20 10 10 1.0
A.1 Data Source and Metadata Collection
Complementary to section 3.1, Table 8 provides a detailed overview of the cyber threat databases we utilize, along with the corresponding metadata collected for various CTI tasks (❶-❻) and analytical targets.
MITRE-CVE and NVD. Our collected data is primarily centered around CVEs, leveraging the MITRE-CVE [14] and NVD [59] databases, which serve as well-structured repositories of publicly disclosed vulnerabilities. These databases provide foundational information, including CVE identifiers, descriptions, severity scores, affected products, and references to external sources.
Exploit-DB. To further enrich our dataset and enhance contextual understanding, we incorporate additional cyber threat intelligence (CTI) sources, including Exploit-DB [62]. Exploit-DB provides reports on security advisories and threat intelligence insights, complementing the structured data from MITRE-CVE and NVD. Notably, Exploit-DB is also CVE-centric; however, unlike MITRE-CVE and NVD, it lacks structured connections to additional evidence. Therefore, our data collection primarily focuses on extracting raw threat descriptions and attack vector classifications (e.g., "Local") to supplement our dataset with additional contextual details.
Other MITRE Databases. Beyond these primary databases, we integrate metadata from additional sources to enrich our dataset with diverse perspectives on cyber threats. Specifically, we leverage CWE [16] to systematically categorize software weaknesses, providing a structured foundation for understanding vulnerability root causes. CAPEC [17] offers a comprehensive taxonomy of attack patterns, enabling us to analyze how adversaries exploit known weaknesses in real-world scenarios. MITRE ATT&CK [18] serves as a crucial resource for mapping adversary tactics, techniques, and procedures (TTPs), facilitating in-depth attribution and behavioral analysis.
Third-Party References. Beyond structured frameworks, we incorporate third-party references to capture real-time threat intelligence and contextualized security insights. Among them, GitHub commits to serve as a valuable source for tracking vulnerability disclosures, exploit development, and security patch discussions. Apache security reports [5] provide detailed documentation on vulnerabilities and remediation efforts specific to widely used web and enterprise software. Meanwhile, IBM X-Force threat intelligence [39] contributes curated threat reports, real-world attack observations, and insights into evolving threat landscapes.
The collected metadata, detailed in Table 8, enhances the comprehensiveness of our dataset by offering multifaceted insights into vulnerability exploitation patterns, adversarial attack techniques, and effective mitigation strategies. By integrating these diverse sources, we enable a richer and more contextualized analysis of cyber threats, supporting improved detection, attribution, and defense mechanisms.
A.2 Threat Corpus Generation Details
Complementary to section 3.1, we detail the prompts used for generating and refining the threat corpus, specifically in Prompt 1 and 2. During output sampling, we adjust the temperature and top-p parameters within a range of 0.2 to 1.0 to balance creativity and determinism. An example of a generated threat report is provided in Example 4. Additionally, we present the statistical overview of the corpus used for pre-training in Table 9.
Prompt 1
Below is the prompt template used for threat corpus generation.
[…]
Prompt 2
Below is the prompt used for threat corpus revision.
Objective: You are an expert cybersecurity analyst and technical writer. Your task is to revise the provided cyber threat report by improving its layout and structure while maintaining all critical details. The goal is to $\cdots$ Instructions: Reorganize Content – Adjust the structure to make the report more logical and easier to follow. Consider using sections such as $\cdots$ Improve Readability – Use concise paragraphs, bullet points, and tables where appropriate. Ensure technical jargon is well-explained. Maintain Accuracy – Preserve all details, technical data, and intelligence without altering their meaning. $\cdots$ Report to Revise: [Insert Cyber Threat Report Here] Output:
Example 4
Below is an example of threat report.
Title: Security Advisory: Cleartext Storage Vulnerability in ChatGPT macOS Application Executive Summary: A security vulnerability has been identified in versions of the ChatGPT application for macOS released before July 5, 2024. The application was found to store user conversations in cleartext, making them accessible to other applications on the system. Additionally, the app operated outside of the macOS sandbox, $\cdots$ Vulnerability Details: CVE Identifier: CVE-2024-40594 Description: The ChatGPT macOS application prior to version 2024-07-05 stored user conversations in an unencrypted format in a location accessible by other applications. Furthermore, the application opted out of the macOS sandbox, which is designed to limit the app’s access to system resources and user data $\cdots$ Affected Products: $\cdots$ Impact: The cleartext storage of conversations poses a privacy risk, as sensitive information could be accessed by other applications or users with sufficient privileges on the same system. The lack of sandboxing further exacerbates the issue by allowing the application broader access to system resources than necessary, potentially exposing additional data $\cdots$ Mitigation: Users are advised to update the $\cdots$
Appendix B Fine-Tuning: Complementary Information
Complementary to section 4.1, we present the statistics of the generated fine-tuning dataset in Table 10. Notably, when generating each data instance using a cascading reasoning layout, not all analytical targets are included due to variations in their availability.
Table 10: Statistics of the fine-tuning dataset with a cascading reasoning layout, where each analytical target appears in varying volumes of data. Additionally, different scales of fine-tuning data are used for different CyLens models to align with their respective capacities and optimization needs.
CTI Task Analytical Target # Sample # Fine-Tuning Samples
1B 8B 70B
Threat Actor 37.6K 0.8K 11.5K 37.6K
TTPs 35.5K
Exploit Path 21.8K
Affected System 37.6K
Attack Infra 28.3K
Attack Impact 17.6K
CVE Identifier 37.6K
CWE Identifier 37.6K
CVE Identifier 22.9K
CVSS Metrics 37.6K
EPSS Records 33.1K
Tool Use 15.3K
Code Patch 8.2K
Methodology 32.4K
Advisory 27.5K
Furthermore, CyLens-1B, CyLens-8B, and CyLens-70B utilize different volumes of fine-tuning data, reflecting their respective model capacities and optimization requirements. For fine-tuning CyLens-1B, we select 0.8K samples that contain all analytical targets. In contrast, for CyLens-8B and CyLens-70B, we use a progressively larger dataset, with CyLens-8B utilizing a partial subset and CyLens-70B leveraging the full dataset, despite the potential absence of certain analytical targets in some instances.
Appendix C Evaluation: Complementary Information
C.1 Evaluation Data
Rationale. Evaluating CTI using both historical cyber threats and zero-day threats provides a comprehensive assessment of their robustness and adaptability. Historical threats serve as a fundamental benchmark, allowing for the validation of defensive mechanisms against well-documented attack patterns, known TTPs. This retrospective analysis ensures that security measures remain effective against previously encountered threats and can detect variations of known exploits. On the other hand, zero-day threats introduce an element of unpredictability, testing the system’s ability to detect and respond to novel attack vectors without predefined signatures or heuristics.
By incorporating both types of threats, evaluations can measure not only the efficiency of existing threat detection models but also the adaptability of security solutions in dynamic and evolving threat landscapes. This dual approach is essential for developing resilient cybersecurity frameworks that balance historical insights with forward-looking threat intelligence.
Table 11: Statistics of evaluation dataset
Historical Zero-Day
# Threats 10,000 5,526
Table 11 presents the volume of our evaluation datasets. We ensure that no threats included in the evaluation appear in the training data.
C.2 Evaluation Metrics
We define the evaluation metrics used for each analytical target, along with the strategies and rationale behind their selection.
    • Threat Actor: Since threat actors are typically unique in each threat report, we use Accuracy to assess whether the threat actor generated by LLMs matches the reference. Additionally, we employ BERT Score to measure the textual similarity between the generated explanation and the reference for threat actor attribution. BERT Score is implemented using the Hugging Face library [^2] https://huggingface.co/spaces/evaluate-metric/bertscore.
    • TTPs: As TTPs are typically presented as a list, we adopt four metrics:
          Precision – Measures the proportion of correctly identified TTPs among those predicted by the LLM.
          Recall – Measures the proportion of actual TTPs that were successfully identified.
          F1 Score – Provides a balanced measure of precision and recall.
          Intersection over Union (IoU) – Quantifies the overlap between the predicted and reference TTP sets.
    • Exploitation Path: Since exploitation paths are usually descriptive observations, we use BERT Score to measure text similarity between the generated and reference descriptions.
    • Affected System & Attack Infrastructure: Similar to TTPs, both affected systems and attack infrastructures are structured as itemized lists. We therefore apply Precision, Recall, F1 Score, and IoU to assess the overlap between the LLM-generated results and the reference lists.
    • Impact: As impact analysis consists of descriptive statements, we use BERT Score to evaluate textual similarity between generated and reference statements.
    • CVE and CWE Detection & Correlation: Given the large-scale CVE candidates and hundreds of CWEs, the detection and correlation tasks prompt LLMs to generate the top-10 most relevant identifiers (CVE or CWE). We use Hit@10 to measure whether at least one correct identifier is present in the top-10 results.
    • Correct Product & Threat Category: To assess whether the generated CVEs correspond to the correct affected product or fall within the correct threat category, we compute Accuracy as the evaluation metric.
    • CVSS Metrics: Since CVSS metrics have a limited number of categorical values (e.g., base severity levels), we use classification accuracy to evaluate the correctness of the predicted severity levels.
    • EPSS Trend Prediction: Since the EPSS score is a numerical probability (e.g., a 5.5% chance of exploitation within a given timeframe), we calculate the Root Mean Square Error (RMSE) to measure the deviation between the LLM-predicted score and the ground truth.
    • Tool Use & Advisory Identification: Both of these tasks involve selecting a single correct item, so we evaluate performance using Accuracy to measure whether the LLM can identify the correct tool or advisory.
    • Code Patch & Mitigation: Since both tasks involve generating textual recommendations, we use BERT Score to assess the similarity between the LLM-generated output and the reference.